A weight suspended from free end of a vertically hanging spring produces an extension of 3 cm . The spring is cut into two parts so that the lenght of the longer part is 2/3 of the original lenght . If the same weight is now suspended from longer part of the spring the extension produced is 

Grade:10

1 Answers

Vikas TU
7 years ago
Weight is same hence, For same mass mg force,
k1x1 = k2x2
k*l = k’*(2l/3)
we get,
k’ =  3k/2
 
FOR natural length l spring,
0.5k3^2 = mg*3
3k/2 = mg................(1)
 
 
For 2l/3 length,
0.5k’x^2 = mgx
we get,
x = 2mg/(3k/2) 
from eq. (1)
we get,
x = 2 cm.